# Graph Linearization Methods for Reasoning on Graphs with Large Language Models

### v1

@@ Abstract
Large language models have evolved to process multiple modalities beyond text, such as images and audio, which motivates us to explore how to effectively leverage them for graph machine learning tasks. The key question, therefore, is how to transform graphs into linear sequences of tokens—a process we term “graph linearization”—so that LLMs can handle graphs naturally. We consider that graphs should be linearized meaningfully to reflect certain properties of natural language text, such as local dependency and global alignment, in order to ease contemporary LLMs, trained on trillions of textual tokens, better understand graphs. To achieve this, we developed several graph linearization methods based on graph centrality, degeneracy, and node relabeling schemes. We then investigated their effect on LLM performance in graph reasoning tasks. Experimental results on synthetic graphs demonstrate the effectiveness of our methods compared to random linearization baselines. Our work introduces novel graph representations suitable for LLMs, contributing to the potential integration of graph machine learning with the trend of multi-modal processing using a unified transformer model.
@@
Christos Xypolopoulos1,3, Guokan Shang2, Xiao Fei1, Giannis Nikolentzos4, Hadi Abdine2, Iakovos Evdaimon1, Michail Chatzianastasis1, Giorgos Stamou3, Michalis Vazirgiannis1,2 1Ecole Polytechnique, 2MBZUAI, 3NTUA, 4University of Peloponnese
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally.
@@ 1 Introduction
Transformer-based large pre-trained models have revolutionized machine learning research, demonstrating unprecedented performance across diverse data modalities and even a mixture of modalities, including image, audio, and text domains (Xu et al., 2023; Yin et al., 2023). Particularly, Large Language Models (LLMs) have shown promising results on arithmetic, symbolic, and logical reasoning tasks (Hendrycks et al., 2020). Despite their success, the adaptation for processing graphs—an ubiquitous data structure that encapsulates rich structural and relational information—remains a comparably emerging and underdeveloped research direction, even if it has recently been gaining attention (Ye et al., 2023; Fatemi et al., 2023; Wang et al., 2024). This asymmetry is due in large part to the inherent challenge of representing graphs as sequential tokens, in a manner conducive to the language modeling objectives typical of transformers, a challenge not encountered when dealing with the other modalities. This unique problem has encouraged us to investigate a key question: How can we represent graphs as linear sequences of tokens for transformers in a suitable way? We refer to this research endeavor as Graph Linearization.
@@
Existing methods of using LLMs for graph machine learning tasks, such as graph reasoning and graph generation, represent entire graphs as either raw edge lists or natural language descriptions that adhere to adjacency matrices without any special treatment (Fatemi et al., 2023; Wang et al., 2024; Yao et al., 2024). For example, a natural language description of the star graph $S_{3}$ might be: “An undirected graph with nodes $a$, $b$, $c$, and $d$. Node $b$ is connected to $a$. Node $c$ is connected to $b$. Node $b$ is connected to $d$.”, its equivalent edge list representation is: “[($b$, $a$), ($c$, $b$), ($b$, $d$)]”. Either of the linearized representations is then appended with a task-specific question to form an LLM query prompt, e.g., “Is there a cycle in this graph?”. Other studies focus solely on node-level tasks (Zhao et al., 2023; Ye et al., 2023), centering the linearization around an ego-subgraph for a target node (Hamilton et al., 2017), where the neighboring graph structure and node features up to $k$-hop are described in the prompt. However, there is a lack of studies investigating how to maintain the integrity of graph structures while efficiently transforming them into sequences suitable for LLMs.
Figure 1: An example of different Graph Linearization methods applied to the motifs’ shape classification task.
@@
Our research addresses this limitation. By relying on edge list representations as exemplified above, we study the performance impact on LLMs of various methods for ordering the edges in the list and renaming interchangeable node labels, as shown in Figure 1. We argue that the linearization of graphs should be conducted in a meaningful way, capturing properties similar to those found in natural language such as local dependency and global alignment, will benefit contemporary LLMs by enhancing their ability to understand graphs, as they are trained on trillions of textual tokens. We define local dependency as the ability to predict the next (missing) token based on the previous (surrounding) context, as seen when a sequence of tokens naturally progresses in text, analogous to the distributional hypothesis of language (Joos, 1950; Harris, 1954; Firth, 1957). Global alignment involves starting sequences from tokens with similar characteristics, aligning the sequences to mimic how text samples begin or end with common words like “The” or “In conclusion” respectively. Addressing this question helps identify LLM-suitable representations for graphs that potentially align with natural languages, unlocking new insights and applications in fields where graphs naturally represent data. Graph linearization paves the way for extending transformer models’ capabilities to graph data and unifying various graph tasks across domains, serving as a fundamental step toward building successful large graph models. This also facilitates the integration of graph learning with the multi-modal processing trend and the development of cohesive AI systems using a unified transformer model.
@@
In this work, we developed multiple general graph linearization methods that leverage graph centrality, degeneracy, and node relabeling techniques to achieve the aforementioned natural language properties. We conducted a series of inference experiments on three basic graph reasoning tasks: node counting, max degree calculation, and motifs’ shape classification, which form the foundation for a deeper understanding of graph structures. Experimental results using the Llama 3 Instruct (Dubey et al., 2024) model on the GraphWave (Donnat et al., 2018) synthetic dataset demonstrate the effectiveness of our methods compared to random linearization baseline.
@@ 2.2 LLMs for graph reasoning
Several studies have evaluated LLMs for graph reasoning at both the node and graph levels. For example, NLGraph (Wang et al., 2024) covers eight graph reasoning tasks of varying complexity, ranging from simple tasks like connectivity and shortest path to complex problems like maximum flow and simulating graph neural networks. This work also proposes two graph-specific prompting methods that achieve notable performance improvements. GraphQA (Fatemi et al., 2023) focuses on relatively simple tasks to measure the performance of pre-trained LLMs in edge existence, node degree, node count, edge count, connected nodes, and cycle checks. It shows that larger models generally perform better on graph reasoning tasks, with graphs generated synthetically using various graph generators. Chen et al. (2023) focuses on node classification on text-attributed graphs. For a target node, this work first uses LLMs to summarize its neighborhood’s textual descriptions, which are then concatenated with the description of the target node to perform classification. Similar works include various studies that explore graph reasoning using different LLMs, prompting techniques, graph tasks, domains, and evaluation approaches (Guo et al., 2023; Zhang et al., 2023a; Hu et al., 2023; Huang et al., 2024; Liu and Wu, 2023; Das et al., 2023; Yuan et al., 2024; Wu et al., 2024; Skianis et al., 2024).
Another line of research criticizes the above approach, arguing that solely using prompt engineering or in-context learning with frozen LLMs hinders achieving top performance in downstream graph tasks. Therefore, instruction-tuning or fine-tuning is necessary. The work of Ye et al. (2023) preliminarily confirms this on the multi-class node classification task. A prompt template is designed to describe both the neighbor graph structure and node features centered around a target node up to the 3-hop level. This template, along with the correct classification answer, is used to fine-tune both encoder-decoder and decoder-only LLMs. Similarly, Zhao et al. (2023) draw inspiration from linguistic syntax trees. For a target node, the work converts its ego-subgraph into a graph syntax tree with branches describing the neighborhood’s “label” and “feature”, which are then encoded as text in the prompt. Results show that instruction-tuning performs much better that in-context learning, and is on par with GNN-based models. Similar conclusions can be observed in recent studies on graph-level reasoning settings (Luo et al., 2024).
@@ 2.3 LLMs for graph generation
To the best of our knowledge, (Yao et al., 2024) was the first to show the preliminary abilities of LLMs in graph generation. The work experimented with graph generation in three settings: 1) Rule-based: generating graphs of basic structure types, given rules describing the desired structures, e.g., trees, cycles, etc.; 2) Distribution-based: generating graphs following a structural type distribution $p$, given a set of example graphs with the same distribution; 3) Property-based: generating molecule structures with specific properties, given example molecules in SMILES format. Results show that LLMs have reasonably good abilities across the three tasks, and advanced prompting techniques do not necessarily lead to better performance.
@@ 2.4 Linearization for specific graphs
The scope of our work. Unlike the above linearization methods limited to specific types of graphs, where linearization can be naturally derived to some extent, we focus on general graphs. Furthermore, unlike previous works using LLMs for reasoning and generation tasks, where edge lists are directly leveraged without special treatment, we introduce various linearization methods for ordering the edges in the list and renaming interchangeable node labels to make them suitable for LLMs. Although our work does not involve experiments with graph generation or instruction tuning, our graph linearization methods are general and applicable to various scenarios. This allows for the effective transformation of graph structures into sequences suitable for language models and has the potential to improve performance in both reasoning and generation tasks, with or without fine-tuning.
@@ 3 Graph linearization methods
Generally speaking, we define graph linearization as the process of representing graphs as linear sequences of tokens. In this work, we aim to identify the linearization approaches that will benefit state-of-the-art LLMs by enhancing their ability to understand graphs. We argue that linearized graphs, represented as sequences of tokens, should capture properties similar to those in natural language, given the fact that LLMs are pre-trained on trillions of textual tokens. Such properties should include local dependency and global alignment.
@@
Local dependency refers to the ability to predict the next (missing) token based on the previous (surrounding) context, within the token sequence of a single linearized graph. This property is analogous to the fundamental distributional hypothesis of language (Joos, 1950; Harris, 1954; Firth, 1957), which states that words that occur in similar contexts tend to have similar meanings (or functions). This hypothesis suggests that given a new word, one should be able to figure out its meaning based on the contexts in which it is used. In fact, the masked and casual language modeling for training encoder-only (Devlin et al., 2019) and decoder-only (Radford et al., 2019) language models, can be seen as instantiations of this hypothesis. Global alignment refers to the alignment across the token sequences of different linearized graphs, so they start or end with tokens sharing similar natures. For example, linearization can be always start from the node with the highest degree, and such nodes are all relabelled as index 0 for different graphs. This property is to reflect the fact that text documents often begin or end with common words like “The” or “In conclusion”.
Figure 2: $k$-core decomposition.
@@
By relying on edge list, which defines a graph in terms of its individual connections, we study the performance impact on LLMs of various methods for ordering the edges in the list and renaming interchangeable node labels, as shown in Figure 1. More specificly, we leverage the advances in graph degeneracy and centrality as detailed in below to meet the local dependency property.
Graph degeneracy (Seidman, 1983). Let $G(V,E)$ be an undirected graph with $n=\left|{V}\right|$ nodes and $m=\left|{E}\right|$ edges. A $k$-core of $G$ is a maximal subgraph of $G$ in which every vertex $v$ has at least degree $k$. As shown in Figure 2, the $k$-core decomposition of $G$ forms a hierarchy of nested subgraphs whose cohesiveness and size respectively increase and decrease with $k$. The higher-level cores can be viewed as a filtered version of the graph that captures main information. The Core Number of a node is the highest order of a core that contains this node.
Graph centrality. A key concept in network analysis to determine the influence or importance of nodes based on the structure of the graph. Degree centrality is one of the simplest measures of node importance Freeman (1978). It is defined as the number of edges incident to a node. For an undirected graph with $n=\left|{V}\right|$ nodes and $m=\left|{E}\right|$ edges, the degree centrality $C_{D}(v)$ of a vertex $v\in V$ is calculated as $C_{D}(v)=\deg(v)=\sum_{u\in V}A_{vu}$ where $A$ is the adjacency matrix of the graph, and $A_{vu}=1$ if there is an edge between vertices $v$ and $u$, and 00 otherwise. PageRank is a more sophisticated centrality measure, originally developed by Brin and Page (1998) for ranking web pages. It extends the concept of degree centrality by considering not only the number of links a node has but also the importance of the nodes linking to it. PageRank, although designed for directed graphs, can also be adapted for undirected ones by treating all edges as bidirectional. The PageRank centrality $C_{PR}(v)$ of a node $v\in V$ is computed iteratively using the formula: $C_{PR}(v)=\frac{1-\alpha}{|V|}+\alpha\sum_{u\in\mathcal{N}(v)}\frac{C_{PR}(u)}{\deg^{+}(u)}$ where $\alpha$ is a damping factor typically set to $0.85$, $\mathcal{N}(v)$ represents the set of nodes linking to $v$, and $\deg^{+}(u)$ is the out-degree of node $u$. PageRank centrality effectively captures the notion that connections from highly-ranked nodes contribute more to the ranking of a given node than connections from low-ranked nodes Page et al. (1999).
Our linearization methods. We order the edges (node pairs) in the edge list based on Core Number, Degree and PageRank as described in above. Specifically, first we rank nodes by their corresponding property. Then, we start with the node having the highest property value and select the edges connected to it, arranging these edges in a random order. This process is repeated for nodes with the next highest property values, continuing in descending order. In the case where two or more nodes share an equal value, the order is selected randomly. After the ordering process has concluded, each edge list constitutes a sequence of tokens following a descending order of node importance. Additionally, we examined the case where we order the edges directly instead of the nodes. To do this, we transformed each graph into its corresponding Linegraph. A linegraph $L(G)$ of a graph $G$ is the graph where each edge of $G$ is replaced by a node, and where two edges of $G$ are connected in $L(G)$ if they are incident in $G$. Then we applied the process described previously using again the core number of each node in $L(G)$.
Finally, we rely on node relabeling to achieve the property of global alignment. For example, the node with index 0 constantly is the one that has the highest core number and so on. This might be beneficial for the LLM to consistently maintain a connection between node indexes and their corresponding importance properties. We evaluate all aforementioned linearization methods with and without this node relabeling strategy.
4 Dataset
@@
We constructed a synthetic graph dataset using GraphWave (Donnat et al., 2018). This graph generator was originally developed for controlled experimentation and evaluation of node embedding techniques and measurement of structural equivalence on graphs with known network motifs. Motifs, in the context of network science, are sub-graphs that repeat themselves either within the same graph or across different graphs. For our analysis, these structurally similar sub-graphs enable us to construct structure-related tasks, that allow us to assess the ability of language models to analyze and infer structural features within these graphs.
@@
The generator operates by constructing sequentially a base graph, that follows either a cycle or chain structure, and then attaching a number of motifs that follow predetermined shapes, to form the final graphs. From the available motif’s shapes, we chose to adopt five of them—cliques, stars, fans, diamonds, and trees—excluding shapes that share structural similarities with the selected ones. To cover a wider variety of structural complexities, we also include all the combinations of two shapes per graph, along with all combinations of three shapes with unique shapes per triplet. Example graphs can be found in Figure 3.

### v3

@@ Abstract
Large language models have evolved to process multiple modalities beyond text, such as images and audio, which motivates us to explore how to effectively leverage them for graph reasoning tasks. The key question, therefore, is how to transform graphs into linear sequences of tokens—a process we term “graph linearization”—so that LLMs can handle graphs naturally. We consider that graphs should be linearized meaningfully to reflect certain properties of natural language text, such as local dependency and global alignment, in order to ease contemporary LLMs, trained on trillions of textual tokens, better understand graphs. To achieve this, we developed several graph linearization methods based on graph centrality and degeneracy. These methods are further enhanced using node relabeling techniques. The experimental results demonstrate the effectiveness of our methods compared to the random linearization baseline. Our work introduces novel graph representations suitable for LLMs, contributing to the potential integration of graph machine learning with the trend of multimodal processing using a unified transformer model.
@@
Christos Xypolopoulos1,3, Guokan Shang2, Xiao Fei1, Giannis Nikolentzos4, Hadi Abdine2, Iakovos Evdaimon1, Michail Chatzianastasis1, Giorgos Stamou3, Michalis Vazirgiannis1,2 1Ecole Polytechnique, 2MBZUAI, 3NTUA, 4University of Peloponnese Correspondence: cxypolop@lix.polytechnique.fr, guokan.shang@mbzuai.ac.ae
@@ 1 Introduction
Transformer-based large pre-trained models have revolutionized machine learning research, demonstrating unprecedented performance across diverse data modalities and even a mixture of modalities, including image, audio, and text domains (Xu et al., 2023; Yin et al., 2023). In particular, large language models (LLMs) have shown promising results in arithmetic, symbolic, and logical reasoning tasks (Hendrycks et al., 2020). Despite their success, the adaptation for processing graphs—an ubiquitous data structure that encapsulates rich structural and relational information—remains a comparably emerging and underdeveloped research direction, even if it has recently been gaining attention (Ye et al., 2023; Fatemi et al., 2023; Wang et al., 2024). This asymmetry is due in large part to the inherent challenge of representing graphs as sequential tokens, in a manner conducive to the language modeling objectives typical of transformers, a challenge not encountered when dealing with the other modalities. This unique problem has encouraged us to investigate a key question: How can we represent graphs as linear sequences of tokens for transformers in a suitable way? We refer to this research endeavor as Graph Linearization.
@@
Existing methods of using LLMs for graph machine learning tasks, such as graph reasoning and graph generation, represent entire graphs as either raw edge lists or natural language descriptions that adhere to adjacency matrices without any special treatment (Fatemi et al., 2023; Wang et al., 2024; Yao et al., 2024). For example, a natural language description of the star graph $S$ might be: “An undirected graph with nodes $a$, $b$, $c$, and $d$. Node $b$ is connected to $a$. Node $c$ is connected to $b$. Node $b$ is connected to $d$.”, its equivalent edge list representation is: “[($b$, $a$), ($c$, $b$), ($b$, $d$)]”. Either of the linearized representations is then appended with a task-specific question to form an LLM query prompt, e.g., “Is there a cycle in this graph?”. Other studies focus solely on node-level tasks (Zhao et al., 2023; Ye et al., 2023), centering the linearization around an ego-subgraph for a target node (Hamilton et al., 2017), where the neighboring graph structure and node features up to $k$-hop are described in the prompt. However, there is a lack of studies investigating how to maintain the integrity of graph structures while efficiently transforming them into sequences suitable for LLMs.
Figure 1: An example of degree-based graph linearization method. Given an input graph G, we rank its nodes based on their degree and then explore the edges in that order. The resulting linearized graph is then combined with a task-specific prompt and passed into a LLM.
@@
Our research addresses this limitation. By relying on edge list representations as exemplified above, we study the performance impact on LLMs of various methods to order the edges in the list and rename interchangeable node labels, as shown in Figure 1. We argue that if the linearization of graphs is conducted in a meaningful way, capturing properties similar to those found in natural language such as local dependency and global alignment, it will benefit contemporary LLMs by enhancing their ability to understand graphs, as they are trained on trillions of textual tokens. We define local dependency as the ability to predict the next (missing) token based on the previous (surrounding) context, as seen when a sequence of tokens naturally progresses in text, analogously to the distributional hypothesis of language (Joos, 1950; Harris, 1954; Firth, 1957). Global alignment involves starting sequences from tokens with similar characteristics, aligning the sequences to mimic how text samples begin or end with common words like “The” or “In conclusion” respectively. Addressing this question helps identify LLM-suitable representations for graphs that potentially align with natural languages, unlocking new insights and applications in fields where graphs naturally represent data. Graph linearization paves the way for extending transformer models’ capabilities to graph data and unifying various graph tasks across domains, serving as a fundamental step toward building successful large graph models. This also facilitates the integration of graph learning with the multi-modal processing trend and the development of cohesive AI systems using a unified transformer model.
@@
In this work, we developed multiple general graph linearization methods that leverage graph centrality and degeneracy, enriched by node relabeling techniques to achieve the aforementioned natural language properties. We conducted a series of inference experiments on various graph reasoning tasks, which form the basis for a deeper understanding of graph structures. Experimental results using Llama 3 models (Dubey et al., 2024) on synthetic datasets demonstrate the effectiveness of our methods compared to the random linearization baseline. Our key findings are as follows:
$\bullet$ Edge ordering matters – Structured graph linearization improves the graph reasoning capabilities of LLMs.
$\bullet$ Node labels contribute significantly – Node labels based on graph features further improve performance.
$\bullet$ Task-specific linearizations – Node-based and edge-based linearizations perform better in respective tasks.
@@ 2.2 LLMs for graph reasoning
Several studies have evaluated LLMs for graph reasoning at both the node and graph levels. For example, NLGraph (Wang et al., 2024) covers eight graph reasoning tasks of varying complexity, ranging from simple tasks like connectivity and shortest path to complex problems like maximum flow and simulating graph neural networks. This work also proposes two graph-specific prompting methods that achieve notable performance improvements. GraphQA (Fatemi et al., 2023) focuses on relatively simple tasks to measure the performance of pre-trained LLMs in edge existence, node degree, node count, edge count, connected nodes, and cycle checks. It shows that larger models generally perform better on graph reasoning, with graphs generated synthetically using various graph generators. Similar works include various studies that explore graph reasoning using different LLMs, prompting techniques, graph tasks, domains, and evaluation approaches (Chen et al., 2023; Guo et al., 2023; Zhang et al., 2023a; Hu et al., 2023; Huang et al., 2024; Liu and Wu, 2023; Das et al., 2023; Yuan et al., 2024; Wu et al., 2024; Skianis et al., 2024).
Another line of research criticizes the above approach, arguing that solely using prompt engineering or in-context learning with frozen LLMs hinders achieving top performance in downstream graph tasks. Therefore, instruction-tuning or fine-tuning is necessary. The work of Ye et al. (2023) preliminarily confirms this on the multi-class node classification task. A prompt template is designed to describe both the neighbor graph structure and node features centered around a target node up to the 3-hop level. Similarly, Zhao et al. (2023) draw inspiration from linguistic syntax trees. For a target node, the work converts its ego-subgraph into a graph syntax tree with branches describing the neighborhood’s “label” and “feature”, which are then encoded as text in the prompt. Results show that instruction-tuning performs much better than in-context learning, and is on par with GNN-based models. Similar conclusions can be observed in recent studies on graph-level reasoning settings (Luo et al., 2024). Finally, Perozzi et al. (2024) introduce GraphToken, which trains an encoder to create continuous representations, rather than converting graphs into text tokens.
@@ 2.3 LLMs for graph generation
To the best of our knowledge, Yao et al. (2024) was the first to show the preliminary abilities of LLMs in graph generation. The work experimented with graph generation in three settings: 1) Rule-based: generating graphs of basic structure types, given rules describing the desired structures, e.g., trees, cycles, etc.; 2) Distribution-based: generating graphs following a structural type distribution $p$, given a set of example graphs with the same distribution; 3) Property-based: generating molecule structures with specific properties, given example molecules in SMILES format. Results show that LLMs have reasonably good abilities across the three tasks, and advanced prompting techniques do not necessarily lead to better performance. comment Graph generation is critical for applications like drug discovery but presents greater challenges than graph reasoning. (Yao et al., 2024) provide early insights into LLM-driven graph generation, evaluating rule-based, distribution-based, and property-based approaches. Their findings suggest that while LLMs can generate structurally meaningful graphs, advanced prompting does not always yield better performance, highlighting the need for more robust generation strategies.
@@ 2.4 Linearization for specific graphs
The scope of our work. Unlike the above linearization methods limited to specific types of graphs, where linearization can be naturally derived to some extent, we focus on general graphs. Furthermore, unlike previous works using LLMs for reasoning and generation tasks, where edge lists are directly leveraged without special treatment, we introduce various linearization methods for ordering the edges in the list and renaming interchangeable node labels to make them suitable for LLMs. Although our work involves only graph reasoning experiments, our graph linearization methods are general and applicable to various scenarios. This allows for the effective transformation of graph structures into sequences suitable for language models and has the potential to improve performance in both reasoning and generation tasks, with or without fine-tuning.
@@ 3 Graph Linearization Methods
This section describes the graph linearization approach, emphasizing the use of graph features to enhance graph reasoning with LLMs.
@@
Generally speaking, we define graph linearization as the process of representing graphs as linear sequences of tokens. In this work, we aim to identify the linearization approaches that will benefit LLMs by enhancing their ability to understand graphs. We argue that linearized graphs, represented as sequences of tokens, should capture properties similar to those in natural language, given the fact that LLMs are pre-trained on trillions of textual tokens. Such properties should include local dependency and global alignment.
@@
Local dependency refers to the ability to predict the next (missing) token based on the previous (surrounding) context, within the token sequence of a single linearized graph. This property is analogous to the fundamental distributional hypothesis of language (Joos, 1950; Harris, 1954; Firth, 1957), which states that words that occur in similar contexts tend to have similar meanings (or functions). This hypothesis suggests that given a new word, one should be able to figure out its meaning based on the contexts in which it is used. In fact, the masked and casual language modeling for training encoder-only (Devlin et al., 2019) and decoder-only (Radford et al., 2019) language models, can be seen as instantiations of this hypothesis.
Global alignment refers to the alignment across the token sequences of different linearized graphs, ensuring that the corresponding tokens of both sequences match across their full length. This property takes into account the overall structure of the sequences, reflecting the typical flow of text, where common words like “The” or “In conclusion” are used at the start or end of a sequence, guiding the alignment. For example, linearization should always start from the node with the highest degree, and such nodes are all relabeled as index 0 for different graphs.
By relying on edge list, which defines a graph in terms of its individual connections, we study the performance impact on LLMs of various methods for ordering the edges in the list and renaming interchangeable node labels, as shown in Figure 1. More specifically, we leverage the advances in graph degeneracy and centrality as detailed in below to meet the local dependency property.
Graph Degeneracy (Seidman, 1983). Let $G(V,E)$ be an undirected graph with $n=\left|{V}\right|$ nodes and $m=\left|{E}\right|$ edges. A $k$-core of $G$ is a maximal subgraph of $G$ in which every node $v$ has at least degree $k$. The $k$-core decomposition of $G$ forms a hierarchy of nested subgraphs whose cohesiveness and size increase and decrease, respectively, with $k$. Higher-level cores can be viewed as filtered versions of the graph that capture the most significant structural information. The Core Number of a node is the highest order of a core that contains the node.
Graph Centrality. Graph centrality is a key concept in network analysis used to determine the influence or importance of nodes based on the structure of the graph. We consider two types of centrality, one based on the Degree of nodes and one based on PageRank.
Degree centrality is one of the simplest measures of node importance (Freeman, 1978). It is defined as the number of edges incident to a node, making it a measure of local centrality that reflects the node’s immediate connectivity within the graph. For an undirected graph $G$ with $n=\left|{V}\right|$ nodes and $m=\left|{E}\right|$ edges, the degree centrality $D(v)$ of a vertex $v\in V$ is calculated as $D(v)=\sum_{u\in V}A_{vu}$ where $A$ is the adjacency matrix of the graph, and $A_{vu}=1$ if there is an edge between vertices $v$ and $u$, and 00 otherwise.
PageRank is a more sophisticated centrality measure, originally developed by Brin and Page (1998) for ranking web pages. It extends the concept of degree centrality by considering not only the number of links a node has but also the importance of the nodes linking to it. PageRank effectively captures the notion that connections from highly-ranked nodes contribute more to the ranking of a given node than connections from low-ranked nodes (Page et al., 1999). PageRank, although designed for directed graphs, can also be adapted for undirected ones by treating all edges as bidirectional. The PageRank centrality $PR(v)$ of a node $v\in V$ is computed iteratively using the formula: $PR(v)=\frac{1-\alpha}{|V|}+\alpha\sum_{u\in\mathcal{N}(v)}\frac{PR(u)}{\deg^{+}(u)}$ where $\alpha$ is a damping factor typically set to $0.85$, $\mathcal{N}(v)$ represents the set of nodes linking to $v$, and $\deg^{+}(u)$ is the out-degree of node $u$.
Graph Linearization Implementation. Our approach to capitalizing on the local dependency property involves the following steps. Given a graph $G$, we initially rank the nodes by the centrality and degeneracy measures described previously. Then, we begin exploring the nodes by descending order and list the edges connected to it, arranging them in a random order. Each edge is represented as a node pair. In the case where two or more nodes share an equal value, the order is selected randomly. After the ordering process has concluded, each edge list constitutes a sequence of tokens following a descending order of node importance.
In addition to linearization methods, node relabeling is employed as a means to attempt the attainment of the global alignment property. Specifically, node relabeling introduces an additional step to our procedure. After ranking the nodes, their original labels are replaced with their respective positions in the ranking. Consequently, the node with index 0 corresponds to the one with the highest core number, and so forth. This approach may prove advantageous for the LLM by ensuring a consistent association between node indices and their respective importance properties.
Finally, we conducted experiments in which the edges were ordered directly, rather than the nodes. This allows our linearization to directly capture relationships between edges, which can be essential for understanding complex graph structures. To achieve this, each graph was transformed into its corresponding linegraph representation. A linegraph $L(G)$ of a graph $G$ is the graph where each edge of $G$ is replaced by a node, and where two edges of $G$ are connected in $L(G)$ if they are incident in $G$. Subsequently, the previously described processes were applied directly to $L(G)$.
4 Experimental Setup
In this section, we provide a detailed overview of the experimental setup, including the methodologies, resources, and evaluation frameworks used in our experiments.
4.1 Datasets
@@
To better explore the capabilities of LLMs on graph reasoning tasks, we utilized two synthetic datasets.
@@
GraphWave. First, we constructed a synthetic graph dataset using GraphWave (Donnat et al., 2018). This graph generator was originally developed for controlled experimentation and evaluation of node embedding techniques and measurement of structural equivalence on graphs with known network motifs. Motifs, in the context of network science, are sub-graphs that repeat themselves either within the same graph or across different graphs. For our analysis, these structurally similar sub-graphs enable us to construct structure-related tasks, that allow us to assess the ability of language models to analyze and infer structural features within these graphs.
@@
The generator operates by sequentially constructing a base graph, that follows either a cycle or chain structure, and then attaching a number of motifs that follow predetermined shapes—cliques, stars, fans, diamonds, and trees. To cover a wider variety of structural complexities, we also include all combinations of two shapes, along with all combinations of three shapes with unique shapes per triplet. Example graphs can be found in Figure 2.